Circular No. 8478

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ab IN NGC 4617 S. Nakano, Sumoto, Japan, reports the discovery by Koichi Itagaki

(Teppo-cho, Yamagata, Japan) of an apparent supernova (red mag 16.7) on numerous unfiltered CCD frames taken around Feb. 5.638 UT with a 0.30-m f/7.8 reflector. The new object is located at α = 12^{h}41^{m}05^{s}.18, δ = +50^{o}22′56′′.2 (equinox 2000.0), which is 7′′ west and 40′′ south of the center of NGC 4617. Nothing was visible at this location on CCD frames taken by Itagaki back to 2001 (limiting mag 19.5) or on the Digitized Sky Survey. Nakano adds that K. Kadota (Ageo, Saitama-ken, Japan) has confirmed the presence of a new star (mag 16.8) at position end figures 05^{s}.19, 55′′.8 on a CCD frame taken with a 0.25-m reflector on Feb. 5.771.

SUPERNOVA 2005ac Further to IAUC 8476, J. Burket and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images taken on Feb. 4.50 (at mag 18.0) and 5.56 UT (mag 18.2). SN 2005ac is located at α = 13h39m55s.82, δ = +0^{o}52′15′′.0 (equinox 2000.0), which is 2′′.2 west and 0′′.9 south of the nucleus of a compact galaxy. A KAIT image taken on Jan. 22.51 also showed the new object at mag ∼ 18.5, while an image taken on 2004 June 20.21 showed nothing at this position (limiting mag 19.5).

1999 HF_{1}

P. Pravec, P. Kušnirák and L. Šarounová, Ondřejov Observatory, report that their photometric observations — obtained during 2002 Mar. 31– May 16 — confirm that 1999 HF_{1} is a binary system as suggested by Pravec et al. (2002, Icarus 158, 276). They derive an orbital period of the system of 14.03 hr, the primary’s rotation period of 2.3192 hr, and a secondaryto-primary mean-diameter ratio of 0.23 ± 0.03. Further observations, e.g. during March–July 2005 (orbit on MPO 66084) are desirable.

COMET C/2004 U1 (LINEAR) Total-magnitude estimates (visual unless otherwise noted): 2004 Nov.

22.74 UT, 17.3 (K. Kadota, Ageo, Saitama, Japan, 0.25-m reflector + CCD); Dec. 2.64, 13.8 (Kadota); 5.98, 13.2 (P. Guzik, Krosno, Poland, 0.20-m reflector); 13.04, 13.1 (J. J. Gonzalez, Asturias, Spain, 0.20-m reflector); 2005 Jan. 8.53, 13.4 (S. Yoshida, Ibaraki, Japan, 0.40-m reflector); 14.94, 13.5 (R. J. Bouma, Groningen, The Netherlands, 0.31-m reflector).

2005 February 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green